# MMAO-Bench: MultiModal All in One Benchmark Reveals Compositional Law between Uni-modal and Omni-modal in OmniModels

### v1

@@ 3.3.2 Dataset Compression
Regarding the existing large-scale uni-modal benchmarks, to reduce the evaluation cost of large-scale models, we designed a clustering-guided hierarchical sampling method. This method achieves efficient compression and high evaluation consistency through the following steps:
@@
Cluster-based Stratification: Utilize the K-Means algorithm to categorize questions into k clusters, each representing a "model performance similar" question type (e.g., easy questions, difficult questions, etc.).
@@
Finally, in order to ensure statistical stability, we repeat above steps by employing 5 random splits and 10-fold cross-validation. This method determines the optimal sample size through cost-benefit curve analysis, reducing evaluation costs by over 90% while maintaining accuracy (Figure.8).
@@ 3.4 Multi-Step Open-Ended Questions
Cross-modality problem requires the model to understand among multiple modalities with multi-step reasoning. However, existing evaluation benchmarks primarily focus on isolated problems where questions and answers remain independent of each other. Such a single-step evaluation paradigm cannot effectively assess the ability of a model to understand the complex, cross modality tasks that require multiple logical steps. Besides, most benchmarks adopt multiple-choice questions to verify the model’s performance, which reduces the difficulty of questions (e.g. allows for guessing) and may underestimate the model’s deficiencies in complex tasks. To address these issues, we introduce a new evaluation method, Multi-Step Open-Ended Questions (MO), which breaks down complex questions into multiple interdependent sub-tasks and requires the model to solve them step by step in the form of open-ended text. The construction process and example are shown in Figure.5. The multi-step open-ended question is worth 10 points. During the process of providing a reference answer, scores are assigned to each sub-question based on its difficulty and necessity for answering the final question, which facilitates the evaluation of the model’s accuracy at each step.
Experiments (4.4.1) show that compared with single-step evaluation method (e.g. multiple-choice questions), multi-step open-ended questions can effectively observe the ability decay of models in long-chain reasoning, providing a more realistic difficulty for advanced models with stronger discrimination.
@@
Table 2: Cross-modal benchmark results on multimodal models
Model Visual Audio Omni-MC Omni-MO
MiniCPM-o-2.6 41.85 63.10 26.77 21.92
Qwen-2.5-omni-3B 45.71 65.81 27.26 25.72
Baichuan-omni-1.5 49.25 66.50 30.32 20.48
Ming-lite-omni-1.5 48.68 71.70 32.10 26.08
Qwen-2.5-omni-7B 52.81 68.11 33.23 29.12
Gemini-2.0-Flash 66.80 74.41 46.26 40.96
Gemini-2.5-Flash 76.64 75.20 68.23 62.00
Gemini-2.5-Pro 80.10 78.70 76.61 68.84
Table 3: Cross-modal analysis on model ability system
@@
MiniCPM-o-2.6 25.55 36.11 27.75 38.78 22.94 26.36 26.40 26.77
Qwen-2.5-omni-3B 27.74 50.00 32.37 38.78 27.52 22.09 25.28 27.26
Baichuan-omni-1.5 27.74 47.22 31.79 28.57 35.78 27.52 29.75 30.32
Ming-lite-omni-1.5 32.85 55.56 37.57 30.61 30.28 29.46 29.98 32.10
Qwen-2.5-omni-7B 40.88 52.78 43.35 38.78 30.28 27.13 29.31 33.23
Gemini-2.0-Flash 46.72 61.11 49.71 63.27 35.78 44.19 44.97 46.29
Gemini-2.5-Flash 67.88 77.78 69.94 85.71 41.28 75.19 67.56 68.23
Gemini-2.5-Pro 81.02 83.33 81.50 91.84 55.05 78.68 74.72 76.61
@@ 4.1 Experiment Setting
We evaluate omni-models that support text, visual, and audio inputs simultaneously, including open-source models: Qwen-3-omni(Xu et al., 2025b), Qwen-2.5-omni-3B, Qwen-2.5-omni-7B(Xu et al., 2025a), Baichuan-omni-1.5(Li et al., 2025), MiniCPM-o-2.6(Yao et al., 2024), and Ming-lite-omni-1.5(AI et al., 2025), as well as closed-source models: Gemini-2.5-Pro, Gemini-2.5-Flash, and Gemini-2.0-Flash(Comanici et al., 2025). All model integrations strictly adhere to official implementations. In video processing, each model receives raw video and performs frame sampling according to its own sampling strategy.
In the following sections, we conduct comprehensive experiments on MMAO-Bench and try to answer the following questions:
1. How about the performance of current omni-models and their shortage abilities?
2. What is the relationship of uni-modal ability and omni-modal ability?
3. Could the MMAO-Bench provide an effective "All in One" evaluation for omni-models?
@@ 4.2.1 Overall Analysis
As shown in Table 2, the proprietary Gemini series of models substantially outperform the evaluated open-source counterparts across all benchmarks. Gemini-2.5-Pro consistently leads, establishing a clear state-of-the-art.
When analyzing the cross-modal tasks specifically, we observe distinct patterns. On the Omni-MC (Multiple-choice) benchmark, the performance of smaller open-source models like MiniCPM-o-2.6 (26.77) and Qwen-2.5-omni-3B (27.26) approximates that of random guessing (0.25). In contrast, larger models such as Qwen-2.5-omni-7B (33.23) demonstrate nascent capabilities beyond this baseline. Nevertheless, a significant performance gap remains, with the top open-source model trailing Gemini-2.5-Pro (76.61) by over 43 points.
The Omni-MO (multi-step open-ended) benchmark, which demands more complex reasoning, proves to be even more challenging for all models. Performance universally drops compared to Omni-MC. Most open-source models score poorly, with the best among them, Qwen-2.5-omni-7B, achieving only 29.12. This indicates severe limitations in their ability to perform multi-step cross-modal reasoning. While the Gemini models also see a score reduction, they maintain a commanding lead, with Gemini-2.5-Pro (68.84) showcasing a far more robust, though not perfect, reasoning capability. This highlights that complex, multi-step, cross-modal understanding remains a critical and unresolved challenge, especially for the open-source community.
In terms of perception capabilities, cross-modal alignment is more challenging than cross-modal recognition. Among open-source models, Qwen-2.5-omni-7B has the strongest alignment capability at 40.88, while Ming-lite-omni-1.5 has the strongest recognition capability at 55.56. The Gemini series significantly outperforms other models, and Gemini-2.5-Pro shows no significant weakness in alignment compared to its recognition capability, both exceeding 80.
@@
In terms of reasoning capabilities 3, spatial reasoning is currently the most challenging task, with the best-performing Gemini-2.5-Pro achieving only 55.05. Among open-source models, Baichuan-omni-1.5 performs the best with a score of 35.78. Ming-lite-omni-1.5 demonstrates the best performance in complex reasoning and general reasoning among open-source models.
Overall, perception ability is relatively simpler compared to reasoning ability, and the model gap in reasoning ability is more significant. Taking Qwen-2.5-omni-7B and Gemini-2.5-Pro as examples, they differ by 38.2 points in perception and 45.41 points in reasoning. This demonstrates that reasoning ability is an extension of perception ability.
@@ 4.2.2 Top-tier Analysis
Why Gemini-2.5-Pro performances significantly better than other models? We attempt to provide an analysis and some hypotheses. On one hand, it stems from the leading uni-modal understanding ability. On the other hand, from the perspective of the reasoning process, Gemini possesses audio captioning capabilities, and it can also naturally incorporate audio content as a basis for reasoning. Existing open-source models, due to their smaller size, lack reasoning processes in a multimodal context. Limited reasoning mostly relies on text or images, with few involving specific audio content.
However, does Gemini performance good enough? To answer this question, we invited human experts to conduct a competition.
Finding 1. Gemini-2.5-Pro has reached human comparable perception ability in omni-modal understanding, but still has a performance gap on reasoning ability. Compared to human experts, Gemini-2.5-Pro only has 8.3% less performance, which shows humanity comparable intelligence. The comparison of scores for specific ability items can be seen in Figure6. Dived into ability analysis, we find a interesting phenomena that human is more capable on reasoning than perception (81.3% > 74.3%), and perception ability is even lower than model’s. Through interviewing some of the human experts, we find that human may lost some information in video or audio in one hand, and in the other hand human has limited world knowledge than large language models.
@@
To investigate the relationship of uni-modal and omni-modal understanding ability, we conduct regression analysis and ablation experiments. Thanks to the unified ability systems and the high quality of omni-modal samples in MMAO-Bench, we find some interesting observations.
Finding 2. Compositional Law: the performance of omni-modal ability is power-law related to the multiply of uni-modals’ performances. Observing the Table.2, we find that the omni-modal performances are correlation to the multiply of their uni-modals’ performances. In details, the performances of different models are relatively close in audio understanding but show significant differences in vision understanding, and these differences are also pronounced in the omni-modal considering both abilities. Gemini-2.5-Flash and Gemini-2.0-Flash are similar in audio understanding, but for vision modality, Gemini-2.5-Flash outperforms Gemini-2.0-Flash by 9.8, and it is 24.4 stronger in the cross-modal understanding.
Finding 3. The weaker models show "short-board effect", while the stronger models demonstrate emergent ability. We draw the multiply of uni-modal scores as a dot line in the Figure.7. Compared with the cross-modal scores, models around the "random guessing stage" are below the line, while Gemini-2.5-Pro shows a significant gap above the line. Besides, these models demonstrate a power-law tendency that the stronger uni-modal ability can stimulate to much stronger cross-modal ability. This also emphasis that multiple modalities can enhance each other and lead to a higher intelligence.
@@
Figure 7: Relationship between uni-modal ability and cross-modal ability among omni-models.
@@
Audio +Caption +Visual Audio +Caption +Visual Audio +Visual
MiniCPM-o-2.6 26.01 30.06 27.75 24.83 25.28 26.40 25.16 26.77
Qwen-2.5-omni-3B 21.39 29.48 32.37 22.60 24.83 25.28 22.26 27.26
Baichuan-omni-1.5 22.54 30.06 31.79 28.86 31.54 29.75 27.10 30.32
Ming-lite-omni-1.5 26.59 27.17 37.57 25.06 24.38 29.98 25.48 32.10
Qwen-2.5-omni-7B 26.01 33.53 43.35 20.81 25.28 29.31 22.26 33.23
Gemini-2.0-Flash 31.21 53.18 49.71 34.90 51.01 44.97 33.87 46.29
Gemini-2.5-Flash 33.53 68.21 69.94 44.30 70.02 67.56 41.29 68.23
Gemini-2.5-Pro 41.62 78.03 81.50 50.34 73.60 74.72 47.90 76.61

### v2

@@ 3.2.2 Dataset Compression
Regarding the existing large-scale uni-modal benchmarks, to reduce the evaluation cost of large-scale models, we designed a clustering-guided hierarchical sampling (CGHS) method as shown in Figure.2. This method achieves efficient compression and high evaluation consistency through the following steps:
@@
Cluster-based Stratification: Utilize the Kmeans++(Arthur and Vassilvitskii, 2007) algorithm to categorize questions into k clusters, each representing a "model performance similar" question type (e.g., easy questions, difficult questions, etc.).
@@
To ensure statistical stability, we repeat the above steps by using 5 random splits and performing 10-fold cross-validation. This approach identifies the optimal sample size via cost-benefit curve analysis, leading to a reduction in evaluation costs by over 90% while preserving accuracy, as shown in Figure.9.
@@ 3.3 Multi-Step Open-Ended Questions
3.3.1 Question Type Definition
Evaluating the multi-step reasoning capabilities of omni models presents a significant challenge. Real-world problems require models to integrate multi-modal information and execute a sequence of logical steps. However, current automated benchmarks, often relying on Outcome Reward Models (ORMs), typically provide only a binary pass/fail judgment. This approach fails to distinguish between a model that completes 80% of a task and one that fails at 20%, a crucial gap that human evaluators easily perceive. While alternatives like Process Reward Models (PRMs)Lightman et al. (2023) or multi-turn dialoguesReddy et al. (2019) exist, they are hampered by high implementation difficulty, low accuracy, or poor efficiency. Moreover, the prevalence of multiple-choice formats in existing benchmarks is unrepresentative of real-world, open-ended user queries and can mask model deficiencies.
To address these issues, we propose an innovative Multi-Step Open-Ended Question (MO) type, designed for granular and realistic assessment. In the construction of MO dataset, complex problems are first deconstructed by human experts into a series of progressive, interdependent sub-questions. Each sub-question is assigned a score based on its importance, summing to a total of 10 points. During testing, all sub-questions are posed in a single turn, requiring the model to generate a step-by-step open-ended response. This method allows us to precisely quantify how far along a complex reasoning chain a model can proceed, offering a more accurate and insightful measure of its true capabilities. An example is shown in Figure.4.
3.3.2 General Scoring Model
Beside the dataset construction, multi-step open-ended question introduces a new challenge of automated evaluation. To overcome this obstacle, we propose a general scoring model that supports multi-choice question, single-step open-ended question and multi-step open-ended question at the same time. Since the task is to compare the target answer and the predicted answer, we use Qwen3-14B(Yang et al., 2025) as backbone and curate a training dataset as illustrated in Figure.5. One of the critical way to improve accuracy is to group questions into finer types and define appropriate criteria for each types. See details in Appendix.D. Through the human-in-the-loop dataset curation, the scoring model achieves 95% accuracy in out-of-distribution models and benchmarks.
Experiments in Section.4.4.1 show that compared with single-step evaluation method (e.g. multiple-choice questions), multi-step open-ended questions can effectively observe the ability decay of models in long-chain reasoning, providing a more realistic difficulty for advanced models with stronger discrimination.
@@ 3.3.2 General Scoring Model
Figure 5: Training dataset construction and model training pipeline for general scoring model.
Table 2: Omni-modal benchmark results on multimodal models.
Model Audio Visual Omni-MC Omni-MO
Qwen-2.5-Omni-3B 54.40 42.67 27.80 24.76
MiniCPM-O-2.6 56.50 42.27 28.60 23.76
Ming-lite-Omni-1.5 58.30 46.28 28.90 25.48
Baichuan-Omni-1.5 54.10 44.66 29.70 21.04
Qwen-2.5-Omni-7B 60.20 50.68 32.60 27.72
Qwen-3-Omni-30B-A3B 79.40 63.29 42.10 37.08
Gemini-2.0-Flash 70.70 62.76 44.90 38.56
Gemini-2.5-Flash 79.50 69.54 54.30 47.08
Gemini-2.5-Pro 88.40 78.67 70.90 57.32
Table 3: Omni-modal analysis on model ability taxonomy. Cross-modal Recognition refers to the set of other Perception capabilities except Cross-modal Alignment.
@@
Qwen-2.5-Omni-3B 29.84 35.94 33.09 20.65 50.00 20.83 23.98 27.80
MiniCPM-O-2.6 26.70 30.88 28.92 26.62 42.42 26.67 28.40 28.60
Ming-lite-Omni-1.5 28.80 35.94 32.60 24.38 43.94 24.17 26.53 28.90
Baichuan-Omni-1.5 30.89 32.26 31.62 25.87 45.45 28.33 28.57 29.70
Qwen-2.5-Omni-7B 38.22 36.41 37.25 28.11 43.94 26.67 29.59 32.60
Qwen-3-Omni-30B-A3B 53.40 45.16 49.02 38.06 53.03 26.67 37.41 42.10
Gemini-2.0-Flash 43.98 49.77 47.06 45.02 57.58 31.67 43.71 44.90
Gemini-2.5-Flash 56.02 50.69 53.19 61.44 68.18 27.50 55.27 54.30
Gemini-2.5-Pro 74.35 70.05 72.06 75.62 84.85 45.00 70.41 70.90
@@ 4.1 Experiment Setting
We evaluate omni models that support text, visual, and audio inputs simultaneously, including open-source models: Qwen-3-Omni-30B-A3B-Instruct(Xu et al., 2025b), Qwen-2.5-Omni-3B, Qwen-2.5-Omni-7B(Xu et al., 2025a), Baichuan-Omni-1.5(Li et al., 2025), MiniCPM-O-2.6(Yao et al., 2024), and Ming-lite-Omni-1.5(AI et al., 2025), as well as closed-source models: Gemini-2.5-Pro, Gemini-2.5-Flash, and Gemini-2.0-Flash(Comanici et al., 2025). To have a fair comparison between instruct model and thinking model, we adopt similar way in Qwen-3(Xu et al., 2025b) that limits thinking budget to 128 tokens. We apply this restriction to Gemini-2.5-Pro and disable the thinking mode for both Gemini-2.5-Flash and Gemini-2.0-Flash. All the other model integrations strictly adhere to official implementations. In video processing, each model receives raw video and performs frame sampling according to its own sampling strategy.
In the following sections, we conduct comprehensive experiments on UNO-Bench and try to answer the following questions:
  1. 1.
    How about the performance of current omni models and their shortage abilities?
  2. 2.
    What is the relationship of uni-modal ability and omni-modal ability?
  3. 3.
    Could the UNO-Bench provide an effective evaluation for the omni model?
@@ 4.2.1 Overall Analysis
Our main evaluation, summarized in Table.2, reveals a clear performance hierarchy where proprietary models, particularly Gemini-2.5-Pro, establish the state-of-the-art across all benchmarks. Meanwhile, progress within the open-source community is notable, with increased model scale and more training data, exemplified by Qwen-3-Omni-30B, leading to substantial improvements. Furthermore, we observe a strong positive correlation between a model’s performance on the foundational Audio and Visual tasks and its scores on the more demanding Omni benchmarks, suggesting that robust uni-modal perception is a prerequisite for advanced omni-modal understanding.
On the Omni-MC (Multiple-choice) benchmark, which evaluates omni-modal comprehension, smaller open-source models exhibit performance marginally surpassing the random guess baseline (25.00), achieving scores between 27.80 and 29.70. The larger Qwen-3-Omni-30B marks a significant leap, with a score of 42.10 that approaches the performance of entry-level proprietary models like Gemini-2.0-Flash (44.90). Nevertheless, a substantial performance deficit persists when compared to the leading Gemini-2.5-Pro (70.90). This gap highlights the profound difficulty of advanced omni-modal comprehension, even in a multiple-choice format.
The Omni-MO (Multi-Step Open-Ended) benchmark presents a considerably greater challenge, as evidenced by the universal and marked degradation in performance for all models relative to their Omni-MC scores. This degradation reveals a systemic limitation in multi-step omni-modal reasoning. For instance, the leading model, Gemini-2.5-Pro, attained a score of merely 57.32 on this benchmark, reflecting a decline of 13.58 points relative to its performance on the Omni-MC task. In comparison, the highest-scoring open-source model, Qwen-3-Omni-30B, achieved only 37.08 points.
To dissect the models’ core capabilities, we perform a fine-grained analysis based on our proposed ability taxonomy, with detailed results presented in Table.3.
In perception, a notable trend emerges: while smaller models find Recognition easier than Alignment, more powerful models like Qwen-3-Omni-30B-A3B and the Gemini-2.5 series exhibit stronger Alignment capabilities. This suggests that advanced models develop a more sophisticated grasp of inter-modal relationships. Among open-source models, Qwen-3-Omni-30B-A3B achieves the highest perception score (49.02). Gemini-2.5-Pro significantly leads overall, with both its Alignment (74.35) and Recognition (70.05) scores surpassing 70.
@@
In reasoning, Spatial Reasoning is consistently the most challenging task across all models. Even the top-performing Gemini-2.5-Pro only achieves 45.00. Notably, Baichuan-Omni-1.5 demonstrates the best spatial reasoning among open-source models with a score of 28.33. For General and Temporal Reasoning, the new Qwen-3-Omni-30B-A3B establishes itself as the open-source leader.
Overall, reasoning proves to be a more challenging frontier than perception. This is highlighted by the performance gap between the leading proprietary model, Gemini-2.5-Pro, and the best open-source model, Qwen-3-Omni-30B-A3B. The disparity is 23.04 points in Perception (72.06 vs. 49.02) but widens to a more substantial 33.00 points in Reasoning (70.41 vs. 37.41). This indicates that advanced reasoning remains a key differentiator and a primary bottleneck for current multimodal models.
@@ 4.2.2 Top-tier Analysis
Why Gemini-2.5-Pro performances significantly better than other models? We attempt to provide an analysis and some hypotheses. On one hand, it stems from the leading uni-modal understanding ability. On the other hand, regarding the reasoning mechanism, Gemini is equipped with audio captioning functionalities as indicated in the technical report(Comanici et al., 2025), and illustrated in Figure.6. It can also naturally incorporate audio content as a foundation for reasoning. Existing open-source models, due to their smaller size, lack reasoning processes in a multimodal context. Limited reasoning mostly relies on text or images, with few involving specific audio content.
Figure 6: An example of omni-modal evaluation result. Gemini-2.5-Pro displays audio captions to assist in solving the problem.
However, does Gemini performance good enough? To answer this question, we invited human experts to conduct a competition. It is important to note that, unlike the dataset annotators, the invited human experts have not been exposed to the question or the answer.
Finding 1. Gemini-2.5-Pro has reached human comparable perception ability in omni-modal understanding, yet there remains a gap in its reasoning performance. Compared to human experts, Gemini-2.5-Pro only has 8.3% less performance, which shows humanity comparable intelligence. The comparison of scores for specific ability items can be seen in Figure.7. Upon examining ability analysis, we observe an intriguing phenomenon: humans are more proficient in reasoning as opposed to perception (81.3% compared to 74.3%), which contrasts with the model’s performance. By interviewing various human experts, it becomes evident that humans might miss some information in video or audio formats, and their world knowledge is more limited compared to large language models.
@@
To investigate the relationship of uni-modal and omni-modal understanding ability, we conduct regression analysis and ablation experiments. Thanks to the unified ability taxonomy and the high quality of omni-modal samples in UNO-Bench, we find some interesting observations.
Finding 2. Compositional Law: the effectiveness of omni-modal capability is related to the product of the performances of individual modalities by a power-law. Observing the results in Table.2, we identify a strong correlation between a model’s omni-modal performance and its uni-modal capabilities. To formalize this, we derive a Compositional Law from a general functional form by applying two simplifying principles dictated by the omni-modal tasks proposed in our UNO-Bench. Let’s elaborate on the specifics below.
General Model & Task Constraints. We begin by positing that the omni-modal performance $\mathcal{P}_{\text{Omni}}$ is a function of uni-modal performances $\mathcal{P}_{\text{A}}$ and $\mathcal{P}_{\text{V}}$. A general form can be written as:
$\begin{split}\mathcal{P}_{\text{Omni}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})={}&f_{\text{A}}(\mathcal{P}_{\text{A}})+f_{\text{V}}(\mathcal{P}_{\text{V}})\\ &+f_{\text{I}}(\mathcal{P}_{\text{A}},\mathcal{P}_{\text{V}})+b\end{split}$ (1)
where $f_{\text{A}},f_{\text{V}}$ represent modality independent path contributions, $f_{\text{I}}$ the interaction, and $b$ a baseline performance constant (e.g. random guess).
We arrive at the following result through rigorous mathematical derivation, and the detailed derivation process is provided in the Appendix B.
$\mathcal{P}_{\text{Omni}}=C\cdot\mathcal{P}_{\text{A}}^{\alpha}\mathcal{P}_{\text{V}}^{\beta}+b$ (2)
where $C$ is a scaling constant, and exponents $\alpha,\beta$ model the interaction’s elasticity.
We then posit a fusion symmetry assumption: in end-to-end omni models, the fusion mechanism does not inherently favor one modality over anotherXu et al. (2025b); Yao et al. (2024), implying symmetric scaling behavior. This leads to $\alpha=\beta$. Substituting this into Eq. 10, we arrive at the Omni-modal Compositional Law:
$\mathcal{P}_{\text{Omni}}=C\cdot(\mathcal{P}_{\text{A}}\times\mathcal{P}_{\text{V}})^{\alpha}+b$ (3)
where $\alpha$ is the synergistic exponent, $C$ is a scaling coefficient, and $b$ is a baseline bias.
A non-linear regression on data from leading models (Figure.8) yields the precise empirical formula:
$\mathcal{P}_{\text{Omni}}\approx 1.0332\cdot(\mathcal{P}_{\text{A}}\times\mathcal{P}_{\text{V}})^{2.1918}+0.2422$ (4)
This model demonstrates an exceptional fit, with a coefficient of determination ($R^{2}$) of 0.9759. Analysis of the fitted parameters reveals a clear transition from limited gains to emergent capabilities, driven by the super-linear nature of the law.
Figure 8: The Compositional Law of Omni-modal Performance. Observed omni-modal scores (dots) versus the product of their uni-modal scores. The dashed line represents our fitted law (Eq. 4), which shows a remarkable alignment with the empirical data ($R^{2}=0.9759$). The convex, accelerating curve visualizes the Power-law synergy.
Power-law Synergy and Emergent Ability.
The exponent $\alpha\approx 2.19$ is the most critical discovery, revealing a powerful Power-law synergy. Because $\alpha>1$, the function is convex, meaning the rate of performance gain accelerates. This explains the transition from a "short-board effect" to an "emergent ability" seen in Figure.8:
  • Limited Gains at Low Performance: For models with weaker uni-modal abilities (e.g., MiniCPM-O), the curve is relatively flat. Small improvements in the product of uni-modal scores yield only marginal gains in omni-modal performance. This can be seen as a "short-board effect", where the system is not yet capable of effectively leveraging the combined inputs.
  • Emergent Ability at High Performance: As uni-modal abilities strengthen (e.g., moving towards Gemini-2.5-Pro), the curve steepens dramatically. The same amount of improvement in the uni-modal product now yields a much larger increase in omni-modal performance. This accelerating return on investment is the quantitative signature of emergence, where stronger foundational skills unlock disproportionately powerful combined capabilities.
Interpreting the Coefficients and Benchmark Coherence.
The other parameters complete the picture. The bias term $b\approx 0.2422$ acts as a performance floor. As uni-modal performances approach zero, the system’s output converges to this value, which is strikingly close to the 0.25 random-guess accuracy of our benchmark. The scaling coefficient $C\approx 1.0332$, being remarkably close to unity, indicates a harmonious and naturally scaled system. We attribute this harmony not only to the models’ intrinsic fusion mechanisms but also to the coherent design of our benchmark itself.
Additional fitted models are presented in Appendix.C. We argue that our proposed model is the most natural and interpretable among them. Interestingly, most models indicate that the visual modality tends to offer greater benefits than the audio modality. This phenomenon may be attributable to the relatively weaker visual capabilities of models at the current stage of development.
It is worth emphasizing that this finding is directly attributed to the deliberate design of UNO-Bench. Specifically, it not only ensures a balanced distribution of capabilities across both uni-modal and omni-modal tasks, but also constructs the majority of questions to demand the joint processing of both modalities for resolution.
@@
Audio + Caption + Visual Audio + Caption + Visual Audio + Caption + Visual
Qwen-2.5-Omni-3B 17.76 29.13 33.09 20.07 21.43 23.98 19.12 24.60 27.80
MiniCPM-O-2.6 29.44 29.61 28.92 27.21 29.93 28.40 28.13 29.80 28.60
Ming-lite-Omni-1.5 26.28 31.07 32.60 23.13 21.43 26.53 24.42 25.40 28.90
Baichuan-Omni-1.5 22.14 32.04 31.62 23.81 26.70 28.57 23.12 28.90 29.70
Qwen-2.5-Omni-7B 22.14 30.10 37.25 20.41 25.34 29.59 21.12 27.30 32.60
Qwen-3-Omni-30B-A3B 27.01 46.84 49.02 18.71 39.63 37.41 22.12 42.60 42.10
Gemini-2.0-Flash 25.55 44.17 47.06 29.76 45.58 43.71 28.03 45.00 44.90
Gemini-2.5-Flash 22.63 49.03 53.19 29.08 53.23 55.27 26.43 51.50 54.30
Gemini-2.5-Pro 37.71 63.83 72.06 42.18 65.99 70.41 40.34 65.10 70.90